OpenStudy (anonymous):

3/6 + 1/5 is the common denominator 30 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so is it 6/30

OpenStudy (anonymous):

wait denominator stays the same through out right ?

OpenStudy (anonymous):

oh no. you have to write each fraction as an equivalent fraction with 30 as the denominator

OpenStudy (anonymous):

\[\frac{3}{6}=\frac{15}{30}\] \[\frac{1}{5}=\frac{6}{30}\] now add

OpenStudy (curry):

its 7 / 10